OpenStudy (anonymous):

(12m + 7)(4m + 12) Help Please!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

were you ever taught FOIL?

OpenStudy (anonymous):

im in 7th grade

OpenStudy (anonymous):

Distribute. (12m+7)(4m+12) = 12m(4m+12) + 7(4m+12)

OpenStudy (anonymous):

okay i was just asking. heres a trick for multiplying polynomials. First: Multiply the first term in each polynomial Outer: Mulitply the outer terms Inner: multiply the inner terms Last: Multiply the last terms

OpenStudy (anonymous):

You'll learn foil either next year or your first year in High school.

OpenStudy (idealist10):

(12m+7)(4m+12) 48m^2+28m+144m+84 =48m^2+172m+84 Answer: 48m^2+172m+84

OpenStudy (dan815):

know basic rules of multiplication

OpenStudy (anonymous):

thanks for the help

OpenStudy (idealist10):

You're welcome.

OpenStudy (dan815):

y(a+b)=ya+yb now if y=c+d plugging it back in you get (c+d)(a)+(c+d)(b) =ca+da+cb+cd

OpenStudy (dan815):

(12m + 7)(4m + 12) = (12m + 7)(4m) + (12m + 7)(12) --> continue simplifying